OpenStudy (anonymous):

Please help ;-; Solve the equation. |k + 6| = 3 a. {–3, 9} b. {–3, 3} c. {–9, –3} d. {all real numbers greater than or equal to –9 and less than or equal to –3}

OpenStudy (anonymous):

\[|k + 6| = 3 \] solve two separate equations for \(k\) a) \(k+6=3\) b) \(k+6=-3\)

OpenStudy (anonymous):

both are solved the same subtract 6

OpenStudy (anonymous):

thank you c:

OpenStudy (anonymous):

yw